OpenStudy (anonymous):

2s+t=r;t Solve each formula in terms of the given variable. Can anybody help me with this its math homework and ive got help with 3 but i have 12 in all!! Please and thank you!

OpenStudy (anonymous):

Just get the t by itself... Thus first you would subtract 2s from both sides and it would be like.. 2s-2s+t = r-2s, thus the 2s's on the left would cancel so you would get... t = r-2s